OpenStudy (anonymous):

Where does the graph of y = 2x2 + x − 15 cross the x-axis?

OpenStudy (anonymous):

is that 2x^2 +x -15 you mean @alamojoe9 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so..its a parabola...Let y = 0 and solve for x...then once you get your two factors, you plug in those numbers into your parabola and that would give you the y for both points...then thats your answer

OpenStudy (anonymous):

factor your expression and set to zero: 2x^2 + x -15 = 0 (2x - 5)(x + 3) = 0 solve for x and that's where your graph crosses the x axis.